OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? Answer -5 1/5 5 -1/5

OpenStudy (anonymous):

The formula for slope (m) between any two points is m = Y2 - Y1 / X2 - X1

OpenStudy (anonymous):

You can chose which of the two points you wish to designate Y or X but you have to be consistent

OpenStudy (amistre64):

missing some (..) for proper notation :) but yes

OpenStudy (anonymous):

Use the slope formula for two points \((x_1,y_1)\) and \((x_2,y_2)\): \[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

Yeah How do you do the fraction thing with this equation type?

OpenStudy (anonymous):

*[ \frac{numerator}{denominator} *] Replace * with \

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

so its 1/5?

OpenStudy (anonymous):

no I got -1/5

OpenStudy (anonymous):

try again using the formula

OpenStudy (anonymous):

\[\frac{4-5}{6-1} = \]

OpenStudy (amistre64):

i tend to subtract points first: (6, 4) -(1, 5) ------ 5,-1 which i use to determine slope with

OpenStudy (anonymous):

You can't switch the order of the points for x1 and x2 unless you also change the order for y1 and y2. The have to be the same relative to each other.

OpenStudy (amistre64):

my old eyes tend to get lost in y1s and y2s so I stack them up to keep them straight ;)

OpenStudy (anonymous):

\[ \frac{x}{y} \] sorry just trying it out

OpenStudy (amistre64):

\[slope=\frac{\triangle y}{\triangle x}:)\]

OpenStudy (anonymous):

haha I did it unconsciously but thanks for noticing that I was just trying the fraction type polpak showed me:)

OpenStudy (anonymous):

I always type out my latex. I find it faster than trying to use the editor. Plus you can do a number of cooler things than what the editor has options for.

OpenStudy (amistre64):

it becomes a little disorientating trying to remember not only how to math, but type, and format into latex all at the same time ..

OpenStudy (anonymous):

\[|a| = \cases{ \begin{array}{cc} a &\text{ if } a \ge 0 \\ -a &\text { if } a \lt 0 \end{array} }\] Whee!